2/5 (x-2)=X+4
I got 2x/5-4/5=x+4 and I'm not sure what to do next because of the fractions

Found 2 solutions by scott8148, edjones:
Answer by scott8148(6628)

if you don't like fractions, clear them in the first step

multiplying by 5 gives 2(x-2)=5x+20

2x-4=5x+20 ... -3x=24 ... x=-8

Answer by edjones(8007)

the first thing you should do is get rid of the fraction, if you can. In this case you can without breaking any rules.
2/5(x-2)=x+4
2(x-2)=5(x+4) multiply 5 times each side.
2x-4=5x+20
-24=3x subtract 2x+20 from each side.
x= -8 divide each side by 3
Check:
2/5(-8-2)=-8+4
2/5(-10)=-4
-4=-4
Ed